The Mars Pathfinder Mission
Status Reports - Third Week

18 July 1997, 10:15 a.m. PDT

The Pathfinder lander and Sojourner rover concluded their 14th successful day of operations on the surface of Mars today, JPL engineers reported.

Highlights of the scientific data returned overnight include a series of images that show Sojourner approaching the white rock Scooby Doo and deploying the rover's X-ray spectrometer instrument to study the rock's surface. Imaging data from the lander's camera also included a high-resolution view of the northernmost of the Twin Peaks seen on the horizon from the landing site.

Pathfinder engineers said all subsystems on both the lander and rover performed flawlessly and that no resets of the lander's computer were detected.

On this Martian Day, Sol 14, Earth rise was at 6:07 p.m. PDT yesterday, sunrise was at 9:15 p.m. PDT, Earth set was at 7:46 a.m. PDT and sunset was at 10:12 a.m. PDT. The day's total data return from the Mars station was 58 megabits.

19 July 1997, 10:00 a.m. PDT

Last night's receipt of scientific data from Mars Pathfinder was delayed until tonight due to minor ground station problems that interfered with capturing all of Pathfinder's radio transmissions, mission engineers said today.

A short downlink opportunity and a problem with ground station computers combined to prevent most of Pathfinder's scientific data from being received last night. But engineering data from the rover and lander show that both remain in excellent health as they completed the first day of their third week on the surface of Mars.

"All the telemetry from the lander and rover continue to show that we have two very healthy spacecraft", said project manager Brian Muirhead. "We successfully completed the rover's seven-day prime mission and have finished the first week of its extended mission, and we are half-way through the lander's 30-day prime mission. Everything looks good for continued operations with outstanding science return from both lander and rover", he said.

Last night's scheduled science data return will be retransmitted during the next Mars day, Sol 16, which begins tonight. Engineers also plan to send a new software patch to remove the software bug that had caused the lander's computer to reset itself earlier in the mission. The next downlink session is scheduled to include images of the Martian moon Phobos, along with observations of early morning fog, measurements of the rock Scoobie Doo and images of various features around the lander.

Mission engineers said that overnight, Sojourner had successfully executed commands to move its wheels to scrape off the top layer of dust from the rock Scoobie Doo. The rover's spectrometer was to have then repositioned its sensor to measure the newly revealed surface of the rock. The extended sensor head, however, apparently overshot the edge of the rock and did not make contact. Engineers will analyze data on the position of the rover and its spectrometer and plan to reposition the instrument tonight.

On this Martian Day, Sol 15, Earth rise was at 6:07 p.m. PDT yesterday, sunrise was at 9:55 p.m. PDT, Earth set was at 8:25 a.m. PDT and sunset was at 10:51 a.m. PDT. The day's total data return from the Mars station was 2 megabits.

20 July 1997, 1:00 p.m. PDT

The Pathfinder mission operations team commanded the lander early this morning and did obtain a carrier signal over the high-gain antenna starting at 3:14 a.m. PDT for the normal period of about 66 minutes, but the signal strength was below expected levels and no scientific data was received.

"This told us that the spacecraft was basically healthy but that there was a problem with the telecommunications link", said Project Manager Brian Muirhead. A later attempt to communicate with the lander through its high-gain antenna from 7:03 to 7:27 a.m. PDT was not successful.

"The flight team is assessing the possible causes of the communication problem", said Muirhead. "This morning's problem may be related to some extent to configuration problems between the spacecraft and the Deep Space Network, but more data is needed to fully assess the problem. We are trying to troubleshoot a problem with very little information", he said.

The flight team is preparing sequences for a low-gain antenna communications session for about midnight tonight (July 20, Pacific Time). A communications session with the high-gain antenna is planned for about 3:30 a.m. PDT tomorrow, July 21.

"Since we have only limited windows to communicate with the spacecraft, we must wait patiently for our next opportunity", Muirhead said. "We will go through the usual steps that have worked for us before, and then we will get to the bottom of the problem as we have before". The telecommunications problem is not thought to be related to the reset problem previously experienced by the lander's computer.

The rover remains safely at the rock called Scooby Doo. Earth will rise over the Sagan Memorial Station at 8:07 p.m. PDT today July 20, and sunrise will be at 11:15 p.m. Earth set is at 9:45 a.m. July 21.

21 July 1997, 10:00 a.m. PDT

The Mars Pathfinder flight team successfully reestablished contact with the Pathfinder lander and rover early this morning, completing several communications sessions using both the low-gain and high-gain antennas.

"What a difference a day makes", said Brian Muirhead, Pathfinder project manager. "The project team has successfully regained full communication capability on both the low-gain and high-gain antennas. The team is extremely pleased with our current status".

"Most of the communications problem experienced over the weekend was associated with ground operations, not with the spacecraft on Mars", Muirhead said. "We'll be working to eliminate the cause of these problems in the coming days, as we return to a more normal mode of operations".

The flight team successfully initiated its first low-gain communications session of the Martian day at 10:38 p.m. Pacific Daylight Time on July 20, then began a second low-gain session at 1:36 a.m. July 21. Both sessions were returning data at the low data rate of 40 bits per second. At 3:22 a.m. PDT, the team conducted a third, brief low-gain session at a slightly higher data rate of 150 bits per second.

"All sessions worked perfectly, and we gained all of the basic engineering and telemetry data that had been stored onboard", Muirhead reported. "We verified that all spacecraft subsystems were healthy".

At 4:50 a.m. PDT, the team conducted a brief high-gain antenna session to make sure the high-gain antenna was pointed at Earth. A full high-gain antenna session at 8,200 bits per second was later performed beginning at 6:43 a.m. PDT. The team acquired all data on lander and rover health and completed acquisition of all of the spacecraft engineering data. They also sent a software update to correct sequences onboard the flight computer which have caused it to automatically reset itself.

Tonight's science activities will include downlinking measurements of a white-colored rock named Scooby Doo, and continuing to acquire data from a full resolution color panoramic photograph of the landing site.

On this Martian day, Sol 17, Earth rose over the newly named Sagan Memorial Station at 8:07 p.m. PDT yesterday July 20. Sunrise was at 11:15 p.m. July 20, and Earth set occurred this morning at 9:45 a.m., July 21.

22 July 1997, 12:00 a.m. PDT

Two-and-a-half weeks after landing in an ancient Martian flood basin known as Ares Vallis, Mars Pathfinder has fulfilled all of its primary science goals and continues to operate nearly flawlessly, the flight team reported at today's press briefing.

More than 300 megabits of data have been returned just in the last week, said Dr. Matthew Golombek, Pathfinder project scientist. The rover continues to follow an aggressive series of maneuvers to study rocks and soils identified by the science teams for their interesting features. In addition, the rover's wheel tracks and soil abrasion experiments are beginning to yield new information about the Martian soil, which appears to be finer than talcum powder.

Worldwide interest in the mission has peaked, with more than 400 million hits reported on the Internet today, said Kirk Goodall, Mars Pathfinder Web engineer. Goodall and David Dubov, Mars Pathfinder webmaster, constructed 20 Pathfinder mirror sites prior to landing day to service the public. The most hits received in a single day -- 46 million -- occurred on July 8, Goodall said, which is more than double the number of hits received in a single day during the 1996 Olympic Games in Atlanta, Georgia.

A communications problem experienced last weekend has been resolved, reported Richard Cook, Mars Pathfinder mission manager. The problem was associated with ground operations, which has been required to reconfigure equipment and software on a daily basis, and the necessity of establishing communications links only during the short periods of time each day when the lander's transmitter is on.

Scientists are beginning to learn more about the Martian soil by studying the rover's wheel tracks, asking it to perform soil abrasion experiments and measuring the magnetic properties of dust that is being collected by a magnetic instrument on the rover. Dr. Henry Moore, a rover scientist with the U.S. Geological Survey in Menlo Park, CA, likened the Martian soil to a very fine-grained silt that could be found in Nebraska. The Martian particles are less than 50 microns in diameter, which is finer than talcum powder.

Dr. Peter Smith, University of Arizona, who is principal investigator of the lander camera, described more about the Martian landscape, pointing out a shallow riverbed crossing through the landing site, and rocks in the distance that were washed into this outflow channel from the Martian highlands.

Science activities tonight will take the rover through the "cabbage patch", an area of soil in between Scooby Doo and a light-colored rock named Lamb. The rover will conduct a soil experiment, then turn and move toward Lamb. Scientists will take measurements of the dark soil near that rock before moving Sojourner close enough to place its spectrometer against the rock.

On this Martian day, Sol 18, Earth rose over the Sagan Memorial Station at 8:47 p.m. PDT yesterday, July 21. Sunrise was at 11:54 p.m. July 21 and Earth set occurred this morning (July 22) at 10:25 a.m. PDT.

23 July 1997, 1:30 p.m. PDT

The Mars Pathfinder lander and rover continue to operate flawlessly on the surface of Mars, 19 days after landing in an ancient outflow channel called Ares Vallis.

Pathfinder's 1-foot-tall roving geologist -- named Sojourner -- continues to collect data on crustal materials and rocks in the immediate vicinity to provide scientists with new information on the geology of this region. The Pathfinder lander, on the other hand, has become a virtual weather station, using its wind socks, pressure and temperature sensors around-the-clock to profile the pressure, temperature, density and opacity of the Martian atmosphere.

Two downlink sessions were successfully completed by 11 a.m. today, using the 70-meter (230-foot) antenna of NASA's Deep Space Network facility in Madrid, Spain, reported David Gruel, Mars Pathfinder flight director, for Sol 19. The flight team retrieved a total of 45 megabits of data overnight, most of which was imaging data from the ongoing science experiments.

"The lander and rover are in excellent health and continue to operate flawlessly", Gruel said. "Meteorological data are being gathered around the clock".

First on Sojourner's list of activities tonight is a wheel abrasion experiment, in which the 10.5-kilogram (23-pound) vehicle will turn and dig some of its wheels into the fine Martian sand to measure material properties of the surface. Next the rover will position its alpha proton X-ray spectrometer face-down in the soil next to a rock called "Lamb" and make measurements of the soil's chemical composition.

On this Martian day, Sol 19, Earth rose over the Sagan Memorial Station at 9:30 p.m. PDT yesterday, July 22. Sunrise was at 12:30 a.m. July 23, and Earth set occurred at 11:04 a.m. PDT today.

24 July 1997, 2:30 p.m. PDT

All communications sessions between the Pathfinder lander and rover were successfully completed today, one day short of the mission's three-week anniversary on the surface of Mars.

Sol 20 began when the Earth rose over Mars' horizon at 10:30 p.m. Pacific Daylight Time last night (July 23), enabling the flight team to initiate communications with the spacecraft. The Sun later rose at 1:15 a.m. PDT this morning, supplying the lander and rover with the energy needed to carry out specific tasks.

Communications were carried out using the 70-meter (230-foot) antenna of NASA's Deep Space Network facility in Madrid, Spain. Forty-seven megabits of data during two downlink sessions were returned on Sol 20.

The data indicated that both the lander and rover remain in excellent health and are continuing to operate masterfully. Flight Director Dave Gruel reported that no further flight software resets have occurred since the team sent modified flight software three Sols, or days, ago.

Today's data included numerous images taken for ongoing science experiments. The Imager for Mars Pathfinder (IMP) also completed another section of the 12-color super panorama image of the landing site, then imaged the rover to add to an ongoing "rover movie" that is being assembled. IMP took a final, end-of-the-day photo of Sojourner following completion of its activities.

Sojourner traveled a total of 7/10ths of a meter (2.3 feet) today and performed another soil mechanics experiment that involved staging a "wheely". The last of its activities was to lower the alpha proton X-ray spectrometer onto the soil near the rock named Lamb. Presently, because it is night on Mars, the rover is powered down and using only its battery to operate the spectrometer and gather data on the Martian soil near Lamb. That data will be transmitted to Earth via the lander during the next Martian day, Sol 21, which begins when Earth rises over Mars tonight at 8:48 p.m. PDT.

Activities for Sol 21 will include another rover soil mechanics test, some more autonomous driving and repositioning of Sojourner's spectrometer against the side of Lamb in preparation for data-gathering the following night.

The lander's meteorological experiment reported highs today of minus 2 degrees Celsius (28 degrees Fahrenheit) and morning low temperatures of minus 73 degrees Celsius (minus 99 degrees Fahrenheit). The weather detectors also recorded large fluctuations of 3/10ths millibars in total pressure on the surface of Mars.

On this Martian day, Sol 20, the Earth set at 11:45 a.m. PDT, ending spacecraft communications with Earth for the day. The Sun set at 1 p.m. PDT.


For further information, please visit our website at



An audio update on Pathfinder's status can be heard by calling 1-800-391-6654.


Status reports prepared by:
Office of the Flight Operations Manager
Mars Pathfinder Project
NASA Jet Propulsion Laboratory
California Institute of Technology
Pasadena, CA 91109


To subscribe to the mpf-status list, send e-mail to

with the following statement in the body of the message:

subscribe mpf-status


EDT = PDT + 3 hrs. = UTC - 4 hrs. Venezuelan Time = UTC - 4:30 hrs.

Updated: July 24 '97

Best seen with MS Internet Explorer.

Forward to the 25 to 31 July Status Reports in ARVAL

Back to the 11 to 17 July Status Reports in ARVAL

Back to the Pathfinder Mission Status Reports in ARVAL